OpenStudy (anonymous):

TEST TOMMOROW, sigma notation in systems of equations help needed dearly. PLS

OpenStudy (kirbykirby):

What's the question

OpenStudy (anonymous):

Its a graph, idk how to put it in, points are (0, 5.4) (1,4.8) (3, 3.5) (5, 2.5)

OpenStudy (anonymous):

Ive been stuck for hours on this stuff

OpenStudy (kirbykirby):

WHat are you supposed to do with those points?

OpenStudy (anonymous):

Find the least squares regression line, y=ax+b and then solve for a and b by solving using the summatic notation

OpenStudy (anonymous):

Im gonna sgnal for help

OpenStudy (anonymous):

@cwrw238

OpenStudy (kirbykirby):

Oh ok so you have the formulas for the regression line parameters right? that being: \[b=\frac{\sum_{i=1}^n(x_i-\bar{x})(y_i-\bar{y})}{\sum_{i=1}^n(x_i-\bar{x})^2} \] and \[ a=\bar{y}-b\bar{x}\] those right?

OpenStudy (anonymous):

Yea, exactly

OpenStudy (kirbykirby):

Ok first find \(\bar{x}\), which is the average of the x-coordinates Then find \(\bar{y}\), which is the average of the y-coordinates and tell me what you get.

OpenStudy (anonymous):

2.25 for x and 15.7 for y

OpenStudy (kirbykirby):

o.o are you sure (for the \(\bar{y}\))? Did you maybe forget a decimal place for one of your y-values when finding that

OpenStudy (anonymous):

3.925 didnt divide

OpenStudy (kirbykirby):

I get 16.2 / 4 = 4.05. But that's probably just a small mistype. The next part, let's focus on the numerator. You have to basically take the first x-value subtract it with \(\bar{x}\), and multiply that with first y-value and subtract \(\bar{y}\). Then you basically write \(+\) and do the same procedure for the 2nd x and y-values. And here \(n=4\) since you have 4 points. So: \[ \sum_{i=1}^4(x_i-\bar{x})(y_i-\bar{y})=(0-2.25)(5.4-4.05)+(1-2.25)(4.8-4.05)\\ ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~+(3-2.25)(3.5-4.05)+(5-2.25)(2.5-4.05) \]

OpenStudy (kirbykirby):

So think of it this way: Do everything you see after the sigma notation using 1 value, then write + and skip to the next value, until you go through your whole list

OpenStudy (kirbykirby):

Does it make sense so far?

OpenStudy (anonymous):

Y= -.58x+54 ? is that the answer

OpenStudy (kirbykirby):

I get -0.586x + 5.369

OpenStudy (kirbykirby):

\(a=\bar{y}-b\bar{x}=4.05-(-0.586)(2.25)=5.369\)